The Earth absorbs energy emitted from the sun. If the sun emits energy at a rate such that the emitted power is 3.9x1026W how much power is incident on the earth? (Hint, you may need to look up the Earth-Sun distance and radius of the Earth.)
Answer #1

Here

Power emitted , P = 3.9 *10^26 W

distance between earth , d = 1.496 *10^11 m

intensity at the location of earth , I = P/(4pi * d^2)

I = 3.9 *10^26/(4*pi * (1.496 *10^11)^2)

I = 1386.72W/m^2

Power incident on earth = I * pi * r^2

Power incident on earth = 1386.72 * pi * (6.371 *10^6)^2

Power incident on earth = 1.77 *10^17 W

the Power incident on earth is 1.77 *10^17 W